OpenStudy (anonymous):

Solve 11.32^(x – 1) = 15.7

OpenStudy (y2o2):

\[(x-1) = \log_{11.32}(15.7) \] \[x = \log_{11.32}(15.7)+1 \approx 2.134 \]

OpenStudy (anonymous):

sorrryyy i wrote the problem wrong! :S. It supposed to be: Solve 11.3^(2x – 1) = 15.7

OpenStudy (y2o2):

\[(2x-1) = \log_{11.32}(15.7) \\ 2x = \log_{11.32}(15.7)+1 \]

OpenStudy (y2o2):

find the value of 2x then divide it by 2 to get the value of x

OpenStudy (anonymous):

i am not sure how to do that ... Please give me more imformation

OpenStudy (anonymous):

hellppp